OpenStudy (anonymous):

I've got a tough one for you guys. The table below shows four systems of equations: System 1 System 2 System 3 System 4 4x - 5y = 2 4x -5y = 2 4x - 5y = 2 4x - 5y = 2 3x - y = 8 3x - 8y = 4 13x - 8y = 26 10x + 3y = 15 Which pair of systems will have the same solution? (4 points) A. System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to two times the second equation in system 1. B. System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to two times the second equation in system 2. C. System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to three times the second equation in system 2. D. System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to three times the second equation in system 1.

OpenStudy (anonymous):

@sourwing Can you help me after you're done there?

OpenStudy (astrophysics):

system 1. x = 38/11, y = 26/11

OpenStudy (astrophysics):

system 2. x = -4/17, y = -10/17

OpenStudy (astrophysics):

system 3. x = 38/11, y = 26/11

OpenStudy (astrophysics):

system 4. x = 81/62, y = 20/31 so systems 1 and 3 are the same.

OpenStudy (anonymous):

Thanks Astro!

OpenStudy (astrophysics):

Np

OpenStudy (anonymous):

@Astrophysics

OpenStudy (astrophysics):

It's D, because three times xD

OpenStudy (anonymous):

That makes sense. Thank you twice!

OpenStudy (astrophysics):

You can try it out yourself just multiply first equation by 3 and add, and you'll get D :)